Todd and Jason are brothers. Todd says I am twice as old as Jason was two years ago. The sum of the brothers ages is 38. How old is each brother
Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

Todd and Jason are brothers.
Todd says I am twice as old as Jason was two years ago.
t = 2(j-2)
t = 2j - 4
The sum of the brothers ages is 38.
t + j = 38
replace t with (2j-4)
2j - 4 + j = 38
2j + j = 38 + 4
3j = 42
j = 42/3
j = 14 yrs is Jason
:
I'll let you find Todd's age, check them in the 1st equation

Answer by Alan3354(69443)

Todd might be lying.